# DogFit: Domain-guided Fine-tuning for Efficient Transfer Learning of Diffusion Models

### v6

@@ Conclusion
Acknowledgments
This research was supported by the Natural Sciences and Engineering Research Council of Canada, and the Digital Research Alliance of Canada.
@@ Acknowledgments
Additional Methodology Explanations
@@
General Assumptions. We assume that both the pretrained source model $\epsilon_{\theta_{0}}$ and the fine-tuned model $\epsilon_{\theta}$ share the same architecture and noise schedule $(\alpha_{t})$, and are trained to approximate the same family of score functions under a common sampling procedure.
@@
Traditionally, CFG estimates both the class-conditional (Equation 12) and the marginal (Equation 13) noises with the same diffusion model. With enough data in the target domain, the class-conditional noise can be well estimated during fine-tuning via $\epsilon_{\theta}(x_{t}\mid c,\mathcal{D}^{\mathcal{T}})$. However, learning the target noise estimator $\epsilon_{\theta}(x_{t}\mid\mathcal{D}^{\mathcal{T}})$ faces the risk of undertraining due to only being trained on a small portion of the small target dataset. Since the source model is usually trained on large, diverse datasets, it likely captures a good approximation of the general image manifold. Furthermore, it has been shown that diffusion models have generalizability on other data distributions (deja2022analyzing) and can denoise noisy images from domains other than those they have been trained on (meng2021sdedit). Therefore, we can assume that the pre-trained source model captures a general enough prior, giving us the ability to approximate the unconditional score using the pre-trained source model, leading to the formulation of DoG:
@@
which exactly matches the offset used in DoG sampling (zhong2025domain). Our method takes a step further from this and applies the same guidance offset online. We train a diffusion model to directly predict this guided score instead of separately learning Equations 12 and 13 in the form of DoG.
@@
Note on the assumptions: Prior work on CFG has shown that the linearity assumption holds in practice for low to moderate $w$ values (zheng2023characteristic). Furthermore, our use of an SED to sample $w$ during training (see Section 3 Incorporating Controllability) helps maintain accurate estimation of $\epsilon$ in the unconditional setting. Based on this, we conjecture that both assumptions can approximately hold in practice.
@@
Since the training loss is a squared error between two linear functions of $w$, minimizing this loss for all values of $w$ is equivalent to matching the corresponding coefficients of both functions. That is, to minimize the loss over the full support of $w$, the model must satisfy:
@@
$\epsilon_{\theta}(x_{t}\mid c,w,\mathcal{D}^{\mathcal{T}})=\epsilon+(w-1)\cdot\left(\epsilon_{\theta}(x_{t}\mid c,\mathcal{D}^{\mathcal{T}})-\epsilon_{\theta_{0}}(x_{t})\right),$ (21)
@@
This enables DogFit to retain MG’s class-conditional guidance while avoiding out-of-distribution generation by pulling samples toward the core of the target domain manifold.
@@
According to Equations 12 and 13, Equation 24 can be rewritten as:
$\sigma(w-1)\cdot\text{sg}\left(\nabla_{x_{t}}\log p_{\theta}(x_{t}|\mathcal{D^{T}})-\nabla_{x_{t}}\log p_{\theta}(x_{t})\right)$ (25)
@@
Thus, given that MG uses $\epsilon_{\theta}(x_{t}|\mathcal{D^{T}})$ as an approximation to $\sigma_{t}\nabla_{x_{t}}\log p(x_{t}|\mathcal{D^{T}})$, we can express DogFit as applying MG, but with a correction term that pushes towards a target domain classifier’s gradient direction. $\blacksquare$
@@
Guidance Scheduling on $\text{FD}_{\text{DINOV2}}$
We observe that delaying guidance using the late-start strategy improves both $\text{FD}_{\text{DINOV2}}$ and precision, consistent with our FID-based findings. This supports the idea that postponing guidance allows for stronger and more accurate directional signals once the model has sufficiently stabilized (Fig. 9a). However, unlike FID, applying the cut-off strategy—where guidance is limited to later denoising steps—shows little to no improvement in FDDINOV2 across most domains (Fig. 9 (b)). We believe this is due to a fundamental difference in what the two metrics capture: FID benefits from preserving the global structure inherited from the source model, which is often formed early in the denoising process. In contrast, FDDINOV2 focuses more on semantic alignment with the target domain, which may require guidance to be present throughout the generation, not just in the later steps. This distinction further highlights the importance of using multiple metrics when evaluating generative models.
Qualitative Comparison on SiT-XL/2
@@
+ CFG (ho2022classifier) - 0.68 0.76 0.72 0.82 0.63 0.72
+ DoG (zhong2025domain) 0.83 0.77 0.81 0.82 0.88 0.77 0.81
MG (tang2025diffusion) - 0.68 0.76 0.71 0.81 0.62 0.72
@@
+ CFG (ho2022classifier) - 0.83 0.67 0.30 0.48 0.47 0.52
+ DoG (zhong2025domain) 0.82 0.91 0.73 0.39 0.59 0.61 0.65
MG (tang2025diffusion) - 0.84 0.63 0.28 0.51 0.44 0.54
@@
+ CFG (ho2022classifier) - 0.56 0.66 0.76 0.53 0.59 0.62
+ DoG (zhong2025domain) 0.61 0.50 0.65 0.68 0.48 0.52 0.57
MG (tang2025diffusion) - 0.56 0.67 0.75 0.55 0.59 0.624
@@
+ CFG (ho2022classifier) - 0.22 0.59 0.87 0.61 0.84 0.63
+ DoG (zhong2025domain) 0.44 0.21 0.57 0.83 0.64 0.83 0.62
MG (tang2025diffusion) - 0.22 0.60 0.89 0.60 0.84 0.63
@@
Precision and Recall Analysis
Tab.4 presents a comparison of Precision and Recall (kynkaanniemi2019improved) between DogFit and the baselines on the DiT-XL/2 backbone. Precision reflects the fidelity or quality of generated samples, while Recall measures sample diversity. As expected, all guidance methods consistently improve Precision compared to finetuning while reducing Recall, aligning with the general behavior of guided diffusion models. This effect arises because guidance steers sampling toward high-likelihood regions of the conditional distribution, which inherently narrows the output distribution and suppresses diversity. Our method, both with and without guidance control, matches the performance of DoG in terms of both Precision and Recall in most settings. Notably, DogFit achieves the highest or second-highest Precision in both Inception and DINOV2 metrics across all datasets). In terms of Recall, DogFit and DoG slightly underperform CFG and MG, which we believe is due to the additional domain alignment term further reducing the diversity to prevent out-of-domain target generations (Equation 22). While recent works explore diversity-preserving techniques and scheduling mechanisms for guidance during sampling (kynkaanniemi2024applying; sadat2023cads), analogous techniques for training-time guidance have yet to be developed. We leave the exploration of such methods for training-guided models like DogFit as an exciting direction for future work.
Scaling to Large Resolutions
Tab.6 provides a comparison between DogFit and CFG when fine-tuning DiT-XL/2 pretrained at $512{\times}512$ on Food101 (bossard2014food) at $512{\times}512$. Results indicate that DogFit is fully scalable to larger resolutions, broadening its applicability and demonstrating invariance to image scale. For this experiment, we use a batch size of 12 and use the same hyperparameters as in our main experiments.
Robustness to Stronger Domain Shifts
While our target datasets span a range of distribution gaps, we additionally evaluate transfer on the EuroSat dataset (helber2019eurosat), consisting of 27,000 satelite images with 10 classes, which shares very few features with the source domain. This setting probes whether DogFit’s source prior still provides a better signal than that of the fine-tuned target under extreme shift. We note that as the target domain diverges substantially from the source, the advantages of transfer learning from the source generally diminishes relative to training on the target data from scratch. Nevertheless, our results in Tab.7 show that DogFit still outperforms Fine-tuning and CFG in this case, indicating that leveraging the source prior can still improve generation quality even when transfer benefits are attenuated.
Exponential vs. Uniform Scheduling
Tab.7 shows that sampling the guidance weight $\omega$ from a SEDD (Fig. 7) consistently outperforms uniform schedules across all test-time values of $\omega$. In particular, SEDD yields lower FID while uniform sampling over broader ranges destabilizes learning and collapses diversity, with FID inflating as $\omega$ increases. This empirical trend aligns with our linearity assumption in Proposition 1 which leads to the equivalence of DogFit and DoG (zhong2025domain) under small $\omega$ values (zheng2023characteristic). By concentrating the probability mass near $\omega\!\approx\!1$, SEDD keeps optimization within the proposition’s validity region, preserving the expected guidance behavior and allowing the use of larger test-time guidance values during generation.
@@
To further show the broad applicability of our method, we show the efficacy of DogFit in two different text-to-image experiments, Dreambooth subject driven transfer  (ruiz2023dreambooth), and LoRA style transfer  (hu2022lora) to Pokémon. Note that these settings typically require large guidance values of up to $7.5$ for the model to produce visually pleasing images. However, DogFit is more theoretically grounded when exposed to small-guidance values during training, where it locally matches DoG (zhong2025domain) (see Proposition 1). While controllable guidance (conditioning on $w$) is a natural extension allowing for large test-time $\omega$ values, we focus here on fixed guidance during training due to limited time and leave that extension for future work. Despite this conservative setting, DogFit significantly improves unguided fine-tuning with small guidance values ($\omega=[1.65,1.75]$) and approaches the performance of CFG with high guidance value ($\omega=7.5$).
@@
We apply Dreambooth fine-tuning on Stable Diffusion v1.5 (rombach2022high; podell2023sdxl) to new subjects using a handful of instance images with DogFit. In this regime, the fine-tuning prompts are typically limited to templates such as “a [v] photo of {subject}, while at inference we expect the model to respond to diverse prompts from the source domain. This mismatch makes faithful distillation of text-conditioning challenging. Nevertheless, Fig. 12 and Tab. 8 show that DogFit learns controllable alignment with the prompt and significantly outperforms unguided fine-tuning while approaching CFG on DINO, CLIP-I, and CLIP-T, the standard metrics for subject-driven generation (ruiz2023dreambooth).

### v7

@@ Conclusion
Acknowledgments. This research was supported by the Natural Sciences and Engineering Research Council of Canada, and the Digital Research Alliance of Canada.
@@
Appendix A Additional Methodology Explanations
@@
General Assumptions. We assume that both the pretrained source model $\epsilon_{\theta_{0}}$ and the fine-tuned model $\epsilon_{\theta}$ share the same architecture and noise schedule $\alpha_{t}$, and are trained to approximate the same family of score functions under a common sampling procedure.
@@
Traditionally, CFG estimates both the class-conditional (Eq. 12) and the marginal (Eq. 13) noises with the same diffusion model. With enough data in the target domain, the class-conditional noise can be well estimated during fine-tuning via $\epsilon_{\theta}(x_{t}\mid c,\mathcal{D}^{\mathcal{T}})$. However, learning the target noise estimator $\epsilon_{\theta}(x_{t}\mid\mathcal{D}^{\mathcal{T}})$ faces the risk of under-training due to only being trained on a small portion of the tiny target dataset (typically 10%  (Ho and Salimans 2022). Since the source model is usually trained on large, diverse datasets, it likely captures a good approximation of the general image manifold. Furthermore, it has been shown that diffusion models have generalizability on other data distributions (Deja et al. 2022) and can denoise noisy images from domains other than those they have been trained on (Meng et al. 2021). Therefore, we can assume that the pre-trained source model captures a general enough prior, giving us the ability to approximate the unconditional score using the pre-trained source model. This leads us to the formulation of DoG:
@@
which exactly matches the offset used in DoG sampling (Zhong et al. 2025). Our method takes a step further from this and applies the same guidance offset online. We train a diffusion model to directly predict this guided score instead of separately learning Eq. 12 and Eq. 13 in the form of DoG.
@@
- Note on the assumptions: Prior work on CFG has shown that the linearity assumption holds in practice for low to moderate $w$ values (Zheng and Lan 2023). Furthermore, our use of an SEDD to sample $w$ during training (see Sec. 3 Incorporating Controllability) helps maintain accurate estimation of $\epsilon$ in the unconditional setting. Based on this, we conjecture that both assumptions approximately hold true in practice.
@@
Since the training loss is a squared error between two linear functions of $w$, minimizing this loss for all values of $w$ is equivalent to matching the corresponding coefficients of both functions. This means that to minimize the loss over the full support of $w$, the model must satisfy:
@@
$\epsilon_{\theta}(x_{t}\mid c,w,\mathcal{D}^{\mathcal{T}}$ $)=$ (21)
$\epsilon+(w-1)\cdot\left(\epsilon_{\theta}(x_{t}\mid c,\mathcal{D}^{\mathcal{T}})-\epsilon_{\theta_{0}}(x_{t})\right),$
@@
This enables DogFit to retain MG’s class-conditional guidance while avoiding out-of-distribution generation by pulling samples the target domain manifold’s core.
@@
According to Eq. 12 and Eq. 13, Eq. 24 can be rewritten:
$\epsilon$ $^{\prime}_{\text{{DogFit}}}-\epsilon^{\prime}_{\text{MG}}=$ (25)
$-\sigma(w-1)\cdot\text{sg}\left(\nabla_{x_{t}}\log p_{\theta}(x_{t}|\mathcal{D^{T}})-\nabla_{x_{t}}\log p_{\theta}(x_{t})\right).$
@@
Thus, given that MG uses $\epsilon_{\theta}(x_{t}|\mathcal{D^{T}})$ as an approximation of the term $\sigma_{t}\nabla_{x_{t}}\log p(x_{t}|\mathcal{D^{T}})$, we can express DogFit as applying the MG’s direction, but with a correction term that further pushes the sample towards a target domain classifier’s gradient direction. $\blacksquare$
@@ Appendix B Additional Experiments
+ CFG (Ho and Salimans 2022) - 0.68 0.76 0.72 0.82 0.63 0.72
+ DoG (Zhong et al. 2025) 0.83 0.77 0.81 0.82 0.88 0.77 0.81
MG (Tang et al. 2025) - 0.68 0.76 0.71 0.81 0.62 0.72
@@
+ CFG (Ho and Salimans 2022) - 0.83 0.67 0.30 0.48 0.47 0.52
+ DoG (Zhong et al. 2025) 0.82 0.91 0.73 0.39 0.59 0.61 0.65
MG (Tang et al. 2025) - 0.84 0.63 0.28 0.51 0.44 0.54
@@
+ CFG (Ho and Salimans 2022) - 0.56 0.66 0.76 0.53 0.59 0.62
+ DoG (Zhong et al. 2025) 0.61 0.50 0.65 0.68 0.48 0.52 0.57
MG (Tang et al. 2025) - 0.56 0.67 0.75 0.55 0.59 0.624
@@
+ CFG (Ho and Salimans 2022) - 0.22 0.59 0.87 0.61 0.84 0.63
+ DoG (Zhong et al. 2025) 0.44 0.21 0.57 0.83 0.64 0.83 0.62
MG (Tang et al. 2025) - 0.22 0.60 0.89 0.60 0.84 0.63
@@
Guidance Scheduling on $\text{FD}_{\text{DINOV2}}$
We observe that delaying guidance using the late-start strategy improves both $\text{FD}_{\text{DINOV2}}$ and precision, consistent with our FID-based findings. This supports the idea that postponing guidance allows for stronger and more accurate directional signals once the model has sufficiently stabilized (Fig. 9a). However, unlike FID, applying the cut-off strategy (limiting guidance to later denoising steps) shows little to no improvement in FDDINOV2 across most domains (Fig. 9 (b)). We believe this is due to a fundamental difference in what the two metrics capture: FID benefits from preserving the global structure inherited from the source model, which is often formed early in the denoising process. In contrast, FDDINOV2 focuses more on semantic alignment with the target domain, which may require guidance to be present throughout the generation, not just in the later steps. This distinction further highlights the importance of using multiple metrics when evaluating generative models.
Qualitative Comparison on SiT-XL/2
Precision and Recall Analysis
Tab.4 presents a comparison of Precision and Recall (Kynkäänniemi et al. 2019) between DogFit and the baselines on the DiT-XL/2 backbone. Precision reflects the fidelity or quality of generated samples, while Recall measures sample diversity. As expected, all guidance methods consistently improve Precision compared to fine-tuning while reducing Recall, aligning with the general behavior of guided diffusion models. This effect arises because guidance steers sampling toward high-likelihood regions of the conditional distribution, which inherently narrows the output distribution and suppresses diversity. Our method, both with and without guidance control, matches the performance of DoG in terms of both Precision and Recall in most settings. Notably, DogFit achieves the highest or second-highest Precision in both Inception and DINOV2 metrics across all datasets). In terms of Recall, DogFit and DoG slightly under-perform CFG and MG, which we believe is due to the additional domain alignment term further reducing the diversity to prevent out-of-domain target generations (Eq. 22). While recent works explore diversity-preserving techniques and scheduling mechanisms for guidance during sampling (Kynkäänniemi et al. 2024; Sadat et al. 2023), analogous techniques for training-time guidance are yet to be developed. We leave the exploration of such methods for DogFit as an exciting direction for future work.
@@
Scaling to Large Resolutions
Tab.6 provides a comparison between DogFit and CFG when fine-tuning DiT-XL/2 pretrained at $512{\times}512$ on Food101 (Bossard et al. 2014) at $512{\times}512$. Results indicate that DogFit is fully scalable to larger resolutions, broadening its applicability and demonstrating invariance to image scale. For this experiment, we use a batch size of 12 and use the same hyperparameters as in our main experiments.
Robustness to Stronger Domain Shifts
While our target datasets span a range of different distribution gaps, we additionally evaluate transfer on the EuroSat dataset (Helber et al. 2019), consisting of 27,000 satelite images with 10 classes, which shares very few features with the source domain. This setting probes whether DogFit’s source prior still provides a better signal than that of the fine-tuned target under extreme shift. We note that as the target domain diverges substantially from the source, the advantages of transfer learning from the source generally diminish relative to training on the target data from scratch. Nevertheless, our results in Tab.7 show that DogFit still outperforms Fine-tuning and CFG in this case, indicating that leveraging the source prior can still improve generation quality even when transfer benefits are attenuated.
Exponential vs. Uniform Scheduling
Tab.7 shows that sampling the guidance weight $\omega$ from a SEDD (Fig. 7) consistently outperforms uniform schedules across all test-time values of $\omega$. In particular, SEDD yields lower FID while uniform sampling over broader ranges destabilizes learning and collapses diversity, with FID inflating as $\omega$ increases. This empirical trend aligns with our linearity assumption in Prop, 1 which hints at the equivalence of DogFit and DoG (Zhong et al. 2025) under small $\omega$ values (Zheng and Lan 2023). By concentrating the probability mass near $\omega\!\approx\!1$, SEDD keeps optimization within the proposition’s validity region, preserving the expected guidance behavior and allowing the use of larger test-time guidance values during generation.
@@
To further show the broad applicability of our method, we test DogFit in two different text-to-image transfer settings, Dreambooth subject driven transfer  (Ruiz et al. 2023), and LoRA style transfer  (Hu et al. 2022) to Pokémon. Note that these settings typically require large guidance values of up to $7.5$ for the model to produce visually pleasing images. However, DogFit is more theoretically grounded when exposed to small-guidance values during training, where it locally matches DoG (Zhong et al. 2025) (see Prop. 1). While controllable guidance (conditioning on $w$) is a natural extension allowing for large test-time $\omega$ values, we focus here on fixed guidance during training and leave that extension for future work. Despite this conservative setting, DogFit significantly improves unguided fine-tuning with small guidance values ($\omega=[1.65,1.75]$) and approaches the performance of CFG with high guidance value ($\omega=7.5$).
@@
We apply Dreambooth fine-tuning on Stable Diffusion v1.5 (Rombach et al. 2022; Podell et al. 2023) to novel subjects using a handful of instance images with DogFit. In this regime, the fine-tuning prompts are typically limited to templates such as “a [v] photo of {subject}, while at inference we expect the model to respond to diverse prompts from the source domain. This mismatch makes faithful distillation of text-conditioning challenging. Nevertheless, Fig. 12 and Tab. 8 show that DogFit learns controllable alignment with the prompt and significantly outperforms unguided fine-tuning while approaching CFG on DINO, CLIP-I, and CLIP-T, the standard metrics for subject-driven generation (Ruiz et al. 2023).